# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v1

@@ E.3 Extended Kernel Benchmark
In Figure 5 and 12 we measure the GPU memory used by the kernels. For this, we use the PyTorch torch.cuda.max_memory_allocated API to measure the peak memory allocated during one kernel iteration. We make sure that the memory statistics are reset after each iteration and that the PyTorch caches are cleared before the start of each benchmark.
@@
  • mLSTMsig (TFLA xl_chunk): TFLA kernel for the mLSTM with sigmoid input gate. We set the chunk size to 128, but find chunk size 256 to perform equally well in terms of runtime (see Fig. 12 and 5).
@@
Interestingly, in our experiments we find that even though the non-materialization version uses the least GPU memory of all FLA kernels, it is neither faster nor more memory efficient in training than our TFLA mLSTM kernels (see Figure 12). While Simple GLA is slightly faster (within 3 ms or 15%), it uses almost twice the GPU memory compared to our TFLA mLSTM kernels.
@@ F.2 Simplified FLOP Count Summary for the mLSTM
Table 17: Total Simplified FLOP counts for the chunkwise-parallel, fully parallel and recurrent formulation of mLSTMexp and mLSTMsig. All terms denote the FLOP count per head and chunk. We set all factors $F_{\text{OP}}$ to 1.
@@ G.1 Theoretical Runtime
If the computation and memory operations can be overlapped, the total runtime is the maximum of the two, i.e.
@@ G.3 Roofline Analysis
The roofline model is a performance model that combines the arithmetic intensity and the accelerator specifications to determine the upper bound on the peak performance of a kernel (Williams et al., 2009). To do so, the roofline model plots the arithmetic intensity on the x-axis and the attainable performance in FLOPs per second (FLOPs/s) on the y-axis. For each hardware accelerator we plot the roofline, which is the maximum performance of the accelerator for a given arithmetic intensity as
@@ G.4 Runtime-Optimal Chunk Size
The main advantage of TFLA is that we can choose arbitrary chunk sizes $L$ to trade off between FLOPs and memory IO and minimize the runtime (see Figure 5 and 20). The reason for this tradeoff is that the chunk size $L$ affects the number of FLOPs and the memory IO (e.g. for loading and storing the memory cell states) (see Appendix F). In this section, we use this insight and our theoretical runtime model from Appendix G.1 to determine the theoretical runtime-optimal chunk size for the TFLA mLSTMsig forward kernel.
@@
where $\mathcal{I}_{\text{acc}}=\frac{\alpha_{\text{ FLOPs}{{}}}}{\beta_{\text{ Bytes}{{}}}}$ is the accelerator arithmetic intensity from Equation (117). Compared to the FLOP-optimal chunk size $L_{\text{opt,FLOP}}$ (see Equation (111)), the runtime-optimal chunk size $L_{\text{opt,Runtime}}$ additionally depends on the arithmetic intensity $\mathcal{I}_{\text{acc}}$ of our hardware and the precision of our states $\text{bytes}_{Cmn}$. The runtime-optimal chunk size grows proportional to the square root of the head dimension $d_{hv}$ (i.e. $\mathcal{O}(\sqrt{d_{hv}})$) and the accelerator arithmetic intensity $\mathcal{I}_{\text{acc}}$ (i.e. $\mathcal{O}(\sqrt{\mathcal{I}_{\text{acc}}})$). We visualize these trends in Figure 21 and 22.
@@ G.5 FLOP and Theoretical Runtime Analysis Summary
    With the chunk size, we trade-off runtime vs. GPU memory usage. [Figure 5]
@@
    The chunk size $L$ determines the total amount of compute in FLOPs with $L=1$ matching the FLOP count of the recurrent formulation and $L=T$ matching the FLOP count of the parallel formulation. [Figure 15]
@@
    There exists an optimal chunk size $L\in[1,T]$ at which the FLOP count is minimized.
@@
    A larger chunk size increases the arithmetic intensity of TFLA kernels. [Equation (117), Figure 18]
@@
    By choosing the chunk size, we select whether the kernel is memory-bound or compute bound on a specific hardware. [Figure 19, Figure 18]
@@
    The optimal chunk size should be determined based on the total runtime instead of the FLOPs/s. [Figure 20, Figure 17]
@@
    The runtime optimal chunk size grows proportional to the square root of the head dimension and the accelerator intensity. [Figure 21, Figure 22]
@@
    Newer hardware generations require larger chunk sizes to reach peak performance. [Figure 22, Figure 17]

### v2

@@ E.3 Extended Kernel Benchmark
In Figure 6 and 15 we measure the GPU memory used by the kernels. For this, we use the PyTorch torch.cuda.max_memory_allocated API to measure the peak memory allocated during one kernel iteration. We make sure that the memory statistics are reset after each iteration and that the PyTorch caches are cleared before the start of each benchmark.
@@
  • mLSTMsig (TFLA xl_chunk): TFLA kernel for the mLSTM with sigmoid input gate. We set the chunk size to 128, but find chunk size 256 to perform equally well in terms of runtime (see Fig. 15 and 6).
@@
Interestingly, in our experiments we find that even though the non-materialization version uses the least GPU memory of all FLA kernels, it is neither faster nor more memory efficient in training than our TFLA mLSTM kernels (see Figure 15). While Simple GLA is slightly faster (within 3 ms or 15%), it uses almost twice the GPU memory compared to our TFLA mLSTM kernels. The speed of Simple GLA can be partly explained to the fact that it computes less FLOPs (no input-gate) and the fact that in the forward pass the memory cell states are materialized in bfloat16, while TFLA materializes states in float32 for improved stability, which causes twice the memory IO per state.
@@ F.2 Simplified FLOP Count Summary for the mLSTM
Table 17: Total Simplified FLOP counts for the chunkwise-parallel, fully parallel and recurrent formulation of mLSTMexp and mLSTMsig. All terms denote the FLOP count per head for a full sequence of length $T$. We set all factors $F_{\text{OP}}$ to 1.
@@ G.1 Theoretical Runtime
If the computation and memory operations can be completely overlapped, the total runtime is the maximum of the two, i.e.
@@ G.3 Roofline Analysis
The roofline model is a performance model that combines the arithmetic intensity and the accelerator specifications to determine the upper bound on the peak performance of a kernel (Williams et al., 2009). To do so, the roofline model plots the arithmetic intensity of the algorithm on the x-axis and the attainable performance in FLOPs per second (FLOPs/s) on the y-axis. For each hardware accelerator we plot the roofline, which is the maximum performance of the accelerator for a given arithmetic intensity as
@@ G.4 Runtime-Optimal Chunk Size
The main advantage of TFLA is that we can choose arbitrary chunk sizes $L$ to trade off between FLOPs and memory IO and minimize the runtime (see Figure 6 and 23). The reason for this tradeoff is that the chunk size $L$ affects the number of FLOPs and the memory IO (e.g. for loading and storing the memory cell states) (see Appendix F). In this section, we use this insight and our theoretical runtime model from Appendix G.1 to determine the theoretical runtime-optimal chunk size for the TFLA mLSTMsig forward kernel.
@@
where $\mathcal{I}_{\text{acc}}=\frac{\alpha_{\text{ FLOPs}{{}}}}{\beta_{\text{ Bytes}{{}}}}$ is the accelerator arithmetic intensity from Equation (119). Compared to the FLOP-optimal chunk size $L_{\text{opt,FLOP}}$ (see Equation (113)), the runtime-optimal chunk size $L_{\text{opt,Runtime}}$ additionally depends on the arithmetic intensity $\mathcal{I}_{\text{acc}}$ of our hardware and the precision of our states $\text{bytes}_{Cmn}$. The runtime-optimal chunk size grows proportional to the square root of the head dimension $d_{hv}$ (i.e. $\mathcal{O}(\sqrt{d_{hv}})$) and the accelerator arithmetic intensity $\mathcal{I}_{\text{acc}}$ (i.e. $\mathcal{O}(\sqrt{\mathcal{I}_{\text{acc}}})$). We visualize these trends in Figure 24 and 25.
@@ G.5 FLOP and Theoretical Runtime Analysis Summary
    The chunk size $L$ mediates a trade-off between runtime and GPU memory usage. [Figure 6]
@@
    $L$ determines the total compute in FLOPs: $L=1$ matches the recurrent formulation, while $L=T$ matches the parallel one. [Figure 18]
@@
    There exists an optimal chunk size $L\in[1,T]$ that minimized the total FLOP count.
@@
    Increasing $L$ raises the arithmetic intensity of TFLA kernels. [Equation (119), Figure 21]
@@
    The chunk size determines whether the kernel is memory-bound or compute-bound on a given hardware.
    [Figure 22, Figure 21]
@@
    FLOPs/s alone can be misleading; the optimal chunk size should be chosen based on total runtime.
    [Figure 23, Figure 20]
@@
    he runtime-optimal chunk size scales proportionally with the square root of the head dimension and the accelerator’s computational intensity.
    [Figure 24, Figure 25]
@@
    Newer hardware generations require larger chunk sizes to approach peak performance. [Figure 25, Figure 20]